OpenStudy (anonymous):

Can someone give me like 5 math word phrases with variables and the translation to the phrase? Medal + Fan .

OpenStudy (agent0smith):

Phrases like this...? The sum of a number and 3 is equal to the product of the number and 6 Which means (if you call the number x) x + 3 = 6x

OpenStudy (anonymous):

Yes just like that

OpenStudy (agent0smith):

The quotient of the square of a number and 2 is the same as the difference of the number and 2 \[\Large \frac{ x^2 }{ 2 } = x- 2\]

OpenStudy (anonymous):

okaii

OpenStudy (agent0smith):

umm... the square of the difference of a number and 5 is the same as the number increased by 10 \[\Large (x-5)^2 = x + 10\]

OpenStudy (anonymous):

Alriii

OpenStudy (agent0smith):

You can come up with two more by probably just changing some wording and numbers in those... try it and post them and let you know if they're fine.

OpenStudy (anonymous):

Okay , 8 more than 5 times a number 5x + 8 lol simple

OpenStudy (agent0smith):

excellent! :) but you have to still make it equal to something... 8 more than 5 times a number is ...

OpenStudy (anonymous):

is... o.o

OpenStudy (agent0smith):

haha just make something up :) is equal to... something.

OpenStudy (anonymous):

7 lol?

OpenStudy (agent0smith):

Sure :D you could use anything you like. "is equal to 7" "is equal to the quotient of the number and 5" "is equal to the number decreased by 9" etc

OpenStudy (anonymous):

lol the first one.

OpenStudy (agent0smith):

Yeah that's fine :D but i mean you can make it equal to anything, so yours is 5x+8 = 7

OpenStudy (anonymous):

Yeah